A property owner listed his property for​ 160% more than he paid for it. The owner eventually accepted an offer 12.5%
below his asking price and sold the property for ​$191 comma 100. How much did the owner pay for the​ property?

Answer by Theo(13342)

he listed the property for 160% more than he paid for it.

so, if the property cost him x dollars, then he listed it for x + 1.6 * x = 2.6 * x.

he sold it for 12.5% less than the listed price.

this means he sold it for 2.6 * x - .125 * 2.6 * x = 2.275 * x.

since he sold it for 191,100, then 2.275 * x = 191,100.

solve for x to get x = 191,100 / 2.275 = 84,000.

to confirm, work your way back up.

he bought the house for 84,000.

he listed it for 160% more than he paid for it.

that means he listed it for 84,000 plus 1.6 * 84,000 = 218,400.

he eventually sold it for 12.5% less than the list price.

218,400 minus .125 * 218,400 = 191,100.

solution checks out ok.

solution is that the owner paid 84,000 for the property.